Imagine a 1.00 L container which initially contains 0.250 atm of H2O and CO each construct and complete an ICE table and determine the equilibrium concentrations of each gas
Answer #1

Solution:

The equilibrium reaction between CO and H2O using ICE table can be written as,

CO + H2O ==== CO2 + H2.

0.250 ---- 0.250 --------- 0 ---------- 0 atm (Initial)

-X -------- -X ---------------- + X ---------- + X atm (Change)

(0.250 - X)--(0.250 - X) --(X) -------X atm. ( Equilibrium)

Thus, equilbrium constant (Kp) is written as,

Kp = [CO2] [H2] / [CO] [H2O]

Kp = X . X / (0.250-X) ( 0.250-X)

Kp = X^2 / (0.250 - X) (0.250-X)

Therefore,

Equilibrium concentration of H2O, CO, H2 and CO2 are:

[H2O] = (0.250-X)

[CO] = (0.250 - X)

[CO2] = X

[H2] = X